OpenStudy (anonymous):

A portrait studio associate completed sales to 5 customers with sales totals of $200, $75, $150, $125 and $10, respectively. What is the average of the associate’s sales to these 5 customers?

OpenStudy (anonymous):

(200+75+150+125+10)/5 = average average = 560/5 average = 112